OpenStudy (anonymous):

what is the radian measure of the smaller angle formed by the hands of a 7 o clock

OpenStudy (anonymous):

work out the degree and convert that into radians using 180 degrees= pi radians :D

OpenStudy (anonymous):

the clock is divided in 12 equal parts. So each part subtends 360/12 = 30 degrees at the center. Now ,at 7, the 2 hands are separated by 5 such parts. So, angle b/w them will be 5*30=150 degrees.

OpenStudy (anonymous):

Also the other angle will be 360-150=210

OpenStudy (anonymous):

therefore the radians = 150 X (pi/180)

OpenStudy (anonymous):

this is over the smaller angle as you requested